An electric blanket uses 21V to heat up when it is switched on. In 5 mins, the blankets circuit transfers 15kJ of energy to the thermal energy store of its surroundings.
Calculate the resistance of the circuit in the electric blanket
1
Expert's answer
2019-11-21T10:05:22-0500

we find the current strength from the Joule-Lenz law:

"Q=I\cdot U\cdot t"

"I=\frac{Q}{U\cdot t}= \frac{15000}{21\cdot 300}= 2.38 A"

Then, according to Ohm's law, we find resistance:

"R=\frac{U}{I}=\frac{21}{2.38}=8.82 Ohm"